# Magnetic Flux Emergence in Binary Neutron Star Remnants

### v1

@@ V Conclusion
Nevertheless, we also observe that these flux tubes can only overcome the stably stratified profile of the remnant if the magnetic field is significantly more than $10^{17}~{}\textrm{G}$, which is often considered unphysically large. In agreement with Musolino et al. [43], this suggests that the strong magnetically-driven outflows needed for BNS remnants to power relativistic jets may originate in the disk rather than the remnant.
@@ Appendix A Equation of State Slices
where $x=\log{\rho}$. We use 100 profiles of the $z$-axis equally spaced in time from the last $4.7$ ms of the remnant studied in Ref. [11] (${\sim}100~{}\mathrm{ms}$ in duration). For fits of $Y_{l}$, we compute $Y_{l}=Y_{e}+Y_{\nu_{e}}-Y_{\bar{\nu}_{e}}$, where $Y_{e}$, $Y_{\nu_{e}}$, and $Y_{\bar{\nu}_{e}}$ are the electron, electron neutrino, and electron antineutrino fractions, respectively. These fits are shown in the left half of Fig. 11.
@@
Figure 11: (Left) Fits of $T$ and $Y_{l}$ as functions of $\rho$ to the remnant from Ref. [11]. Black lines indicate profiles along the z-axis from the last $4.7$ msof the remnant, the red line marks the fit to this line. (Right) Mass-radius curves for the hot DD2 profile used in this work and the standard cold beta-equilibrium slice of DD2.
@@ Appendix B Relativistic Magnetic Buoyancy in the TOV Spacetime
Note: for this section, we consider geometrized units with $G=c=k_{b}=1$ and adopt the Heaviside-Lorentz convention for the magnetic field; that is, $B^{i}_{\mathrm{HL}}=B^{i}_{\mathrm{CGS}}/\sqrt{4\pi}$. We also use the Einstein summation convention for paired indices, e.g., $a^{\mu}b_{\mu}\equiv\sum_{\mu}a^{\mu}b_{\mu}$.
Consider the relativistic generalization of Eq. 1 [44]:
@@
Eq. 24 is often a weaker constraint than its Newtonian counterpart. A massive remnant may have $m(r)\sim 2~{}\textrm{M}_{\odot}$ near a shear layer located at $r\sim 2~{}\textrm{M}_{\odot}$, suggesting that $(m(r)+4\pi r^{3}P_{0})/(r-2m(r))$ can be more than a factor of $2$ larger than $m(r)/r$. The precise behavior of the term $\Delta e-\Delta P$ depends on the actual EOS, but a sufficient condition to ensure that $\Delta e-\Delta P>\Delta\rho$ is that $d(\rho\epsilon)/dP>1$. This is always satisfied for an ideal gas respecting causality ($1<\Gamma\leq 2$), though it may not always hold for a realistic EOS. Nevertheless, we have found for our cases that the product $(\Delta e-\Delta P)(1+r\pi r^{3}P_{0}/m(r))(1-2m(r)/r)^{-1}>1$.

### v2

@@ V Conclusion
Nevertheless, we also observe that these flux tubes can only overcome the stably stratified profile of the remnant if the magnetic field is significantly more than $10^{17}~\textrm{G}$, which is often considered unphysically large. In agreement with Musolino et al. (2025), this suggests that the strong magnetically-driven outflows needed for BNS remnants to power relativistic jets may originate in the disk rather than the remnant.
@@ Appendix A Equation of State Slices
where $x=\log{\rho}$. We use 100 profiles of the $z$-axis equally spaced in time from the last $4.7$ ms of the remnant studied in Ref. Radice and Bernuzzi (2023) (${\sim}100~\mathrm{ms}$ in duration). For fits of $Y_{l}$, we compute $Y_{l}=Y_{e}+Y_{\nu_{e}}-Y_{\bar{\nu}_{e}}$, where $Y_{e}$, $Y_{\nu_{e}}$, and $Y_{\bar{\nu}_{e}}$ are the electron, electron neutrino, and electron antineutrino fractions, respectively. These fits are shown in the left half of Fig. 12.
@@
Figure 12: (Left) Fits of $T$ and $Y_{l}$ as functions of $\rho$ to the remnant from Ref. Radice and Bernuzzi (2023). Black lines indicate profiles along the z-axis from the last $4.7$ msof the remnant, the red line marks the fit to this line. (Right) Mass-radius curves for the hot DD2 profile used in this work and the standard cold beta-equilibrium slice of DD2.
@@ Appendix B Relativistic Magnetic Buoyancy in the TOV Spacetime
Note: for this section, we consider geometrized units with $G=c=1$ and adopt the Heaviside-Lorentz convention for the magnetic field; that is, $B^{i}_{\mathrm{HL}}=B^{i}_{\mathrm{CGS}}/\sqrt{4\pi}$. We also use the Einstein summation convention for paired indices, e.g., $a^{\mu}b_{\mu}\equiv\sum_{\mu}a^{\mu}b_{\mu}$.
Consider the relativistic generalization of Eq. 1 Lichnerowicz (1994):
@@
Eq. 24 is often a weaker constraint than its Newtonian counterpart. A massive remnant may have $m(r)\sim 2~\textrm{M}_{\odot}$ near a shear layer located at $r\sim 7~G\textrm{M}_{\odot}/c^{2}\approx 10.3~\mathrm{km}$, suggesting that $(m(r)+4\pi r^{3}P_{0})/(r-2m(r))$ can be more than a factor of $2$ larger than $m(r)/r$. The precise behavior of the term $\Delta e-\Delta P$ depends on the actual EOS, but a sufficient condition to ensure that $\Delta e-\Delta P>\Delta\rho$ is that $d(\rho\epsilon)/dP>1$. This is always satisfied for an ideal gas respecting causality ($1<\Gamma\leq 2$), though it may not always hold for a realistic EOS. Nevertheless, we have found for our cases that the product $(\Delta e-\Delta P)(1+r\pi r^{3}P_{0}/m(r))(1-2m(r)/r)^{-1}>1$.